# InvestESG: A multi-agent reinforcement learning benchmark for studying climate investment as a social dilemma

### v1

@@ 3 The InvestESG Environment
State and Observation Space. The environment simulates a partially observable Markov game $\mathcal{M}$ defined over a continuous, multi-dimensional state space. The system state at period $t$ is characterized by the three climate risk parameters $\mathbf{P}_{t}=(P_{t}^{h},P_{t}^{p},P_{t}^{d})$, each company agent’s state vector $\mathbf{S}_{t}^{\mathcal{C}_{i}}=(K_{t}^{\mathcal{C}_{i}},Q_{t}^{\mathcal{C}_{i}},L_{t}^{\mathcal{C}_{i}})$, where $K_{t}^{\mathcal{C}_{i}}$ is the capital level, $Q_{t}^{\mathcal{C}_{i}}$ is the ESG score, and $L_{t}^{\mathcal{C}_{i}}$ is the climate resilience. Each investor agent’s state is represented by their investment portfolio and cash levels $\mathbf{S}_{t}^{\mathcal{I}_{j}}=(H_{t,1}^{\mathcal{I}_{j}},\dots,H_{t,M}^{\mathcal{I}_{j}},C_{t}^{\mathcal{I}_{j}})$, where $H_{t,i}^{\mathcal{I}_{j}}$ represents investor $\mathcal{I}_{j}$’s holdings in company $\mathcal{C}_{i}$, and $C_{t}^{\mathcal{I}_{j}}$ is the investor’s cash level. The full system state at period $t$ is thus $\mathcal{S}_{t}=\left(\mathbf{P}_{t},\{\mathbf{S}_{t}^{\mathcal{C}_{i}}\}_{i=1}^{M},\{\mathbf{S}_{t}^{\mathcal{I}_{j}}\}_{j=1}^{N}\right)$. All company and investor agents share a common observation space, denoted as $\mathcal{O}_{t}$. In the default setting, $\mathcal{O}_{t}=\left(\{\mathbf{S}_{t}^{\mathcal{C}_{i}}\}_{i=1}^{M},\{\mathbf{S}_{t}^{\mathcal{I}_{j}}\}_{j=1}^{N}\right)$. Extensions that incorporate additional observable information, like climate risk, are explored in Section 4.
State Transition. The environment’s state transition $\mathcal{T}$ proceeds as follows. At the beginning of period $t$, investors collect their investment holdings from period $t-1$ and redistribute their capital according to $\mathbf{a}_{t}^{\mathcal{I}_{j}}$. Denote $||a_{t}||_{1}^{\mathcal{I}_{j}}=\sum_{i=1}^{M}a_{t,i}^{\mathcal{I}_{j}}$ as the number of companies investor $\mathcal{I}_{j}$ invests in during period $t$ and let $K_{t}^{\mathcal{I}_{j}}=\sum_{i=1}^{M}H_{t,i}^{\mathcal{I}_{j}}+C_{t}^{\mathcal{I}_{j}}$ represent the total capital of investor $\mathcal{I}_{j}$ at the start of period $t$. Companies reach an interim capital level after returning old investments, $\sum_{j=1}^{N}H_{t,i}^{\mathcal{I}_{j}}$, and receiving new ones, with the investment from investor $\mathcal{I}_{j}$ calculated as $a_{t,i}^{\mathcal{I}_{j}}\frac{K_{t}^{\mathcal{I}_{j}}}{||a_{t}||_{1}^{\mathcal{I}_{j}}}$ or 00 if the investor opts out of investing, as shown in Equation 4.
$K_{t+1,interim}^{\mathcal{C}_{i}}=K_{t}^{\mathcal{C}_{i}}-\sum_{j=1}^{N}H_{t,i}^{\mathcal{I}_{j}}+\sum_{j=1}^{N}a_{t,i}^{\mathcal{I}_{j}}\frac{K_{t}^{\mathcal{I}_{j}}}{||a_{t}||_{1}^{\mathcal{I}_{j}}},\quad\text{for }i=1,\dots,M$ (4)
Companies then make climate-related spending using the interim capital, as described in Equations 5 to 6. Here, $U_{t,m}$ represents the cumulative mitigation spending by all company agents up to period $t$, while $U_{t,r}^{\mathcal{C}_{i}}$ denotes the cumulative resilience spending by company $\mathcal{C}_{i}$.
$U_{t,m}=U_{t-1,m}+\sum_{i=1}^{M}u_{t,m}^{\mathcal{C}_{i}}\times K_{t+1,interim}^{\mathcal{C}_{i}}$ (5)
$U_{t,r}^{\mathcal{C}_{i}}=U_{t-1,r}^{\mathcal{C}_{i}}+u_{t,r}^{\mathcal{C}_{i}}\times K_{t+1,interim}^{\mathcal{C}_{i}}\quad\text{for }i=1,\dots,M$ (6)
While $U_{t,m}$ is then plugged into Equation 1 where system climate risks are updated. Equation 7 states that a company’s climate resilience, $L_{t}^{\mathcal{C}_{i}}$, scales with the proportion of cumulative resilience investment relative to its capital, and that increasing resilience becomes progressively more challenging due to diminishing returns (Pörtner et al., 2022).
$L_{t}^{\mathcal{C}_{i}}=L_{0}^{\mathcal{C}_{i}}\exp{(-\eta^{\mathcal{C}_{i}}\frac{U_{t-1,r}^{\mathcal{C}_{i}}+u_{t,r}^{\mathcal{C}_{i}}K_{t+1,interim}^{\mathcal{C}_{i}}}{K_{t+1,interim}^{\mathcal{C}_{i}}})},\quad\text{for }i=1,\dots,M.$ (7)
At the same time, the occurrence of climate events are simulated according to Equation 2, and companies receive updated ESG scores based on Equation 3. Afterwards, companies’ profit margins, $\rho_{t}^{\mathcal{C}_{i}}$, are computed using Equation 8, factoring in climate-related spendings $u_{t,m}^{\mathcal{C}_{i}}$, $u_{t,g}^{\mathcal{C}_{i}}$, $u_{t,r}^{\mathcal{C}_{i}}$, default economic growth $\gamma$, and losses due to climate events, which are influenced by companies’ climate vulnerability $L_{t}^{\mathcal{C}_{i}}$ and the number of climate events $X_{t}$ as defined in Equation 2:
$\rho_{t}^{\mathcal{C}_{i}}=(1-u_{t,m}^{\mathcal{C}_{i}}-u_{t,g}^{\mathcal{C}_{i}}-u_{t,r}^{\mathcal{C}_{i}})(1+\gamma)(1-X_{t}L_{t}^{\mathcal{C}_{i}})-1,\quad\text{for }i=1,\dots,M.$ (8)
Company capitals $K_{t+1}^{\mathcal{C}_{i}}$, investor holdings $H_{t+1,i}^{\mathcal{I}_{j}}$, and investor cash positions $C_{t+1}^{\mathcal{I}_{j}}$ are updated according to Equations 9 to 11. Company capitals $K_{t+1}^{\mathcal{C}_{i}}$ are updated according to Equation 9 by scaling the interim capital levels by profit margin.
$K_{t+1}^{\mathcal{C}_{i}}=(1+\rho_{t}^{\mathcal{C}_{i}})K_{t+1,interim}^{\mathcal{C}_{i}},\quad\text{for }i=1,\dots,M.$ (9)
Equation 10 adjusts investor holdings based on company profit margins in their portfolios.
$H_{t+1,i}^{\mathcal{I}_{j}}=a_{t,i}^{\mathcal{I}_{j}}(1+\rho_{t}^{\mathcal{C}_{i}})\frac{K_{t}^{\mathcal{I}_{j}}}{||a_{t}||_{1}^{\mathcal{I}_{j}}},\quad\text{for }i=1,\dots,M,j=1,\dots,N.$ (10)
If an investor chooses not to invest, all capital remains as cash, as shown in Equation 11
$C_{t+1}^{\mathcal{I}_{j}}=\begin{cases}0,&\text{if }||a_{t}||_{1}^{\mathcal{I}_{j}}\neq 0\\ K_{t}^{\mathcal{I}_{j}},&\text{if }||a_{t}||_{1}^{\mathcal{I}_{j}}=0\end{cases},\quad\text{for }j=1,\dots,N.$ (11)
@@
Rewards. The single-period reward for company $\mathcal{C}_{i}$ is solely based on its profit margin, given by $r_{t}^{\mathcal{C}_{i}}=K_{t+1}^{\mathcal{C}_{i}}-K_{t+1,interim}^{\mathcal{C}_{i}}$ for $i=1,\dots,M$, reflecting the assumption that companies are profit-driven. The reward for investor $\mathcal{I}_{j}$ is $r_{t}^{\mathcal{I}_{j}}=\frac{K_{t+1}^{\mathcal{I}_{j}}-K_{t}^{\mathcal{I}_{j}}}{K_{t}^{\mathcal{I}_{j}}}+\alpha^{\mathcal{I}_{j}}\frac{\sum_{i=1}^{M}H_{t+1,i}^{\mathcal{I}_{j}}Q_{t+1}^{\mathcal{C}_{i}}}{\sum_{i=1}^{M}K_{t+1}^{\mathcal{I}_{j}}}$ for $j=1,\dots,N$. The first component is the portfolio return ratio, and the second component represents the weighted average ESG score of the investor’s portfolio adjusted by the investor’s ESGpreference, $\alpha^{\mathcal{I}_{j}}$.
@@
Social Outcome Metrics. We evaluate agent performance based on two key social outcome metrics: the final climate risk level, $P_{100}$, defined as $P_{100}=1-(1-P_{100}^{h})(1-P_{100}^{p})(1-P_{100}^{d})$, and the total market wealth at the end of the period, $W_{100}$, defined as $W_{100}=\sum_{i=1}^{M}K_{100}^{\mathcal{C}_{i}}+\sum_{j=1}^{N}K_{100}^{\mathcal{I}_{j}}$.
@@ 3.1 The trade-off for the agents frames climate change as a social dilemma
With companies balancing between short-term costs and long-term benefits, as well as self-interest and social welfare, and investors weighing investment returns against ESG preferences, these trade-offs create a social dilemma within the environment. To illustrate this, we use empirical Schelling diagrams (Hughes et al., 2018), which plot the focal-company payoffs from following either a cooperative or defecting policy, depending on the number of other cooperating company agents in a 5-company, 3-investor environment.
@@
Figure 3: Schelling diagrams demonstrating that the environment constitutes a social dilemma. Subfigure (a) illustrates the selfish scenario, where all three investors consistently prioritize financial returns ($\alpha^{\mathcal{I}_{j}}=0,\text{ for }j=1,2,3$). Here, the payoff for an individual company for defecting (not investing in reducing emissions) is always higher than cooperating (mitigating), so companies will selfishly choose to defect. However, if all companies defect, overall profits will be lower. Subfigure (b) and (c) correspond to two and three infinitely ESG-conscious investors ($\alpha^{\mathcal{I}_{j}}\approx\infty$), respectively. In (b), when few other companies cooperate (corresponding to points 0, and 1 on the x-axis), cooperation yields a higher individual payoff than defection for the focal company. In (c), cooperation outperforms defection in all cases. Therefore, (b-c) demonstrate that investor behavior can change the properties of the environment such that it no longer represents a social dilemma, in that companies may be greedily motivated to mitigate to attract investors. Subfigure (d) and (e) have the same setting as (c) with three ESG-conscious investors. In addition, (d) adds the option for resilience spending to (c). Subfigure (e) adds the option of greenwashing to (c), which once again converts the environment to a social dilemma in which corporations will no longer invest in mitigation.
@@
In our case, cooperation represents altruistic investment in mitigating emissions. We simulate a cooperative policy as one which consistently invests 0.5% of capital in mitigation and 0 in greenwashing or resilience building. In Figure 3(a) to 3(c), a defector policy takes zero action in all of mitigation, greenwashing, and resilience. As shown in Figure 3(a), when investors are solely profit-driven, the payoff for following a defector policy is always higher than the cooperator payoff. However, if all agents fail to cooperate, overall payoffs will be extremely low, showing that the environment represents a social dilemma, which by definition, involves a conflict between immediate self-interest and longer-term collective interests. Figures 3(b) and 3(c) illustrate how the payoffs for cooperation and defection evolve as the proportion of ESG-conscious investors with positive ESG preferences increases. Notably, Figure 3(b) suggests the possibility of a bifurcated equilibrium when investor preferences are mixed. Specifically, climate-friendly companies may attract ESG-conscious investors, while other companies free ride and attract profit-motivated investors. Figure 3(c) shows that when all investors prioritize ESG-friendly firms, cooperation dominates defection in all cases. Therefore, with enough ESG-conscious investors the environment is no longer a social dilemma, and companies are actually greedily motivated to invest in mitigation to attract investors.
@@ 3.2 Multi-agent Reinforcement Learning Baselines
To test how self-interest agents learn to respond to incentives in the environment, we employ a state-of-the-art MARL algorithm based on Independent PPO (IPPO) (De Witt et al., 2020; Yu et al., 2022). Each agent has its own policy parameters, and agents do not share parameters among themselves. This is because we are interested in simulating companies and investors as independent, selfishly motivated agents that specialize in maximizing their own expected reward.  [^4] In simulating climate events as described in Equation equation 2, we use a fixed random seed across training episodes. This approach ensures that, while event occurrence varies with episode-specific climate risks, the underlying randomness remains deterministic. This setup is based on two considerations. First, unlike MARL agents learning from scratch, real-world companies possess a baseline understanding of trade-offs and can reasonably anticipate future risks; a deterministic training environment mirrors this insight. Second, our study aims not to develop agents that solve climate change under all stochastic scenarios, but rather to train agents that grasp the strategic trade-offs faced by companies and investors. A partially deterministic environment makes learning and convergence easier without compromising the study’s structural conclusions. For robustness, the environment setup allows agents to be trained in a fully stochastic environment to ensure generalizability.
By default, 5 companies are modeled with an initial capital of \$10 trillion each, alongside 3 investors, each starting with \$16 trillion. This creates a total beginning market wealth of \$98 trillion, which is comparable to the global stock market cap (WFE Statistics, 2023). We conduct the experiments in such a 5-company, 3-investor environment by default unless specified otherwise. All the following results are averaged result based on 3 runs of different random seeds, with the standard error plotted in the shaded region.
4 Results
@@
Figure 4: Climate risk, total market wealth, and total mitigation amount over the course of training for the 5-company-3-investor case. We compare the status quo scenario with solely profit-driven investors (investors with ESG consciousness level of 0), both with and without the ESG disclosure mandate, to scenarios involving three ESG-conscious investors with ESG consciousness level of $\alpha=0.5$, $\alpha=1$, and $\alpha=10$. These results indicate that merely disclosing ESG scores is insufficient to resolve the social dilemma if investors are not interested in investing in climate-friendly companies. Only when investors have a high level of ESG consciousness—$\alpha=10$, in this case—does the ESG mandate make a difference in increasing the level of mitigation.
The status quo environment presents a social dilemma. For all experiments, our evaluation metric is (1) ending system climate risk, $P_{100}$, and (2) ending market wealth, $W_{100}$. We first model the agents’ behavior under the status quo where there are no ESG disclosure mandates. All the agents are solely profit-motivated and no ESG score is released. We monitor how the evaluation metrics evolve over the course of training in Figure 4. The company agents begin training by acting randomly, leading to more frequent mitigation. Total market wealth and total reward for individual agents is highest in the first 200 episodes, which is the optimal collective outcome. As the training progresses, the company agents learn to not mitigate in order to selfishly maximize their return. Conflict breaks out when agents discover that, in a situation of great climate risk, investing in mitigation is marginally more beneficial than not, which is shown by a steady and slow increase in total market wealth, converging around \$2,000 trillion. However, the profit-motivated strategy for the individual agents achieves a sub-optimal collective outcome, as is characteristic of a social dilemma.
Does ESG score alone incentivize mitigation? As shown in Figure 4, the ending system climate risks for the status quo scenario is around 0.75. To model the effects of ESG disclosure mandate, we add each company’s annual ESG score to agents’ observation space, calculated according to Equation 3. The learning curve shown in Figure 4 reveals a similar level of climate risk, indicating that mandatory ESG disclosure alone does not significantly incentivize emission mitigation when investors care only about profits.
@@
How conscious do investors need to be? Studies have found that ESG disclosure mandates promote climate-friendly efforts driven by societal and stakeholder pressures (see e.g., Cormier and Magnan 1999; Gamerschlag et al. 2011; Fama and French 2007; Friedman and Heinle 2016; Chen et al. 2018; Ioannou and Serafeim 2019), and that investment fund managers are willing to sacrifice financial returns for ESG benefits (Krueger et al., 2021). Therefore, we study the effects of investors with different level of ESG consciousness on the emission mitigation level. As shown in Figure 4, when all investors have a high level of ESG consciousness ($\alpha=10$), we can see a significant difference where investors incentivize companies to invest in mitigation, lowering final climate risk and increasing market wealth. In this case, a few leading mitigating companies attract the majority of ESG-conscious investment while others remain inactive, as shown in Appendix Figure 9(c). However, ESG consciousness levels of 0.5 and 1 (valuing climate investment and profits equally) are insufficient to resolve the dilemma.
What if investors have different levels of ESG-consciousness? Research shows that investors have varying preferences for ESG efforts and respond in different ways (Amel-Zadeh and Serafeim, 2018). In this scenario, we initialized three investors with different levels of ESG-consciousness, with Investor 0 representing the solely profit-seeking investor with ESG-preference set to $\alpha^{\mathcal{I}_{0}}=0$, and Investor 1 and 2 representing highly ESG-conscious investors with $\alpha^{\mathcal{I}_{1}}=\alpha^{\mathcal{I}_{2}}=10$, which matches the scenario presented in the Schelling diagram in Figure 3(b). Figure 5(a) and 5(b) highlight a bifurcation in both investor and company behavior. Profit-driven investor 0 distributes investments evenly across companies, while ESG-focused investors (1 and 2) favor climate-conscious firms, such as Company 0, which prioritizes mitigation. This bifurcation extends to company strategies: Company 0 learns to attract more ESG investment by focusing on mitigation, while others prioritize financial returns at the expense of some investor interest.
@@
Will companies cheat with greenwashing? Building on research that examines the existence and extent of greenwashing (Wu et al., 2020; Marquis et al., 2016; Siano et al., 2017), and the concern that the ESG disclosure policy can backfire with greenwashing (El-Hage, 2021), we explore whether companies adjust their strategies when greenwashing is permitted. Based on the Schelling diagram in Figure 3(d) and Figure 5(b), we predict that investors would be misled by greenwashing, prompting companies to prioritize greenwashing over mitigation due to its lower cost, leading to waste of resources. However, this prediction is not observed when simulating agents with IPPO. Figure 6(b) plots results under varying greenwashing costs, represented by coefficient $\beta$ in Equation 3, where a larger $\beta$ indicates cheaper greenwashing. All investors have an ESG-consciousness level of $\alpha=1$. The results show that regardless of greenwashing cost, companies initially explore greenwashing and mitigation equally but quickly abandon greenwashing. This likely occurs because, in the early episodes, investors have not yet linked ESG scores to their investment strategies, so do not respond to greenwashing efforts with increased investment. Instead, greenwashing presents an immediate cost for companies, which does not pay off with reduced climate risk, so they quickly learn to avoid it. In contrast, even without immediate investor rewards, companies may recognize the long-term benefits of mitigation (which leads to higher collective returns), and be incentivized to continue those efforts at first, even if they later learn to defect. This mirrors reality, where if investors are slow to adjust their strategies, companies may abandon greenwashing early but persist in low levels of mitigation for either its long-term climate benefits or the foresight of regulations and societal pressure that will eventually nudge them towards sustainable operation anyway. This is consistent with some empirical literature stating that despite the emergence of some greenwashing, ESG disclosure mandates overall encourage mitigation (Fiechter et al., 2022).
@@
Can providing additional information about climate risk affect behavior? In this scenario, we want to observe the effect of having additional climate-related information available on mitigation behavior. Therefore, we provide the climate event probability and climate event occurrences as additional information in the observation space for both companies and investors. In the default 5-company-3-investor scenario, having additional climate-related information reduces the ending system climate risks, as shown in Figures 7(a). Figure 7(b) show similar effects of having additional climate-related information in the scenario where investors are not present. This result suggests that climate-related information helps both companies and investors make better environmental decisions. Even in the absence of ESG-focused investors, educating companies about the overall system improves outcomes. This aligns with literature showing that raising awareness encourages positive corporate responses (Delmas and Toffel, 2008; Bowen, 2000).

### v2

@@ 4 The InvestESG Environment
State and Observation Space. The environment simulates a partially observable Markov game $\mathcal{M}$ defined over a continuous, multi-dimensional state space. The system state at period $t$ is characterized by the climate risk parameter $P_{t}$, each company agent’s state vector $\mathbf{S}_{t}^{\mathcal{C}_{i}}=(K_{t}^{\mathcal{C}_{i}},Q_{t}^{\mathcal{C}_{i}},L_{t}^{\mathcal{C}_{i}})$, where $K_{t}^{\mathcal{C}_{i}}$ is the capital level, $Q_{t}^{\mathcal{C}_{i}}$ is the ESG score, and $L_{t}^{\mathcal{C}_{i}}$ is the climate resilience. Each investor agent’s state is represented by their investment portfolio and cash levels $\mathbf{S}_{t}^{\mathcal{I}_{j}}=(H_{t,1}^{\mathcal{I}_{j}},\dots,H_{t,M}^{\mathcal{I}_{j}},C_{t}^{\mathcal{I}_{j}})$, where $H_{t,i}^{\mathcal{I}_{j}}$ represents investor $\mathcal{I}_{j}$’s holdings in company $\mathcal{C}_{i}$, and $C_{t}^{\mathcal{I}_{j}}$ is the investor’s cash level. The full system state at period $t$ is thus $\mathcal{S}_{t}=\left(P_{t},\{\mathbf{S}_{t}^{\mathcal{C}_{i}}\}_{i=1}^{M},\{\mathbf{S}_{t}^{\mathcal{I}_{j}}\}_{j=1}^{N}\right)$. All company and investor agents share a common observation space, denoted as $\mathcal{O}_{t}$. In the default setting, $\mathcal{O}_{t}=\left(\{\mathbf{S}_{t}^{\mathcal{C}_{i}}\}_{i=1}^{M},\{\mathbf{S}_{t}^{\mathcal{I}_{j}}\}_{j=1}^{N}\right)$. Extensions that incorporate additional observable information, like climate risk, are explored in Section 5.
State Transition. The environment’s state transition $\mathcal{T}$ proceeds as follows. At the beginning of period $t$, investors collect their investment holdings from period $t-1$ and redistribute their capital according to $\mathbf{a}_{t}^{\mathcal{I}_{j}}$. If an investor chooses not to invest, all capital remains as cash. Companies then make climate-related spending, after which system climate risk and companies’ inidividual resilience are updated. Meanwhile, companies receive updated ESG scores, and the occurrence of climate events are simulated. Afterwards, companies’ profit margins, $\rho_{t}^{\mathcal{C}_{i}}$, are computed as a function of their climate-related spending $\mathbf{u}_{t}^{\mathcal{C}_{i}}$, default economic growth $\gamma$, and losses due to climate events, $L_{t}^{\mathcal{C}_{i}}$, and the number of climate events $X_{t}$. Finally, company capitals are updated, so are investor holdings and cash positions based on the profit margins of their portfolios.
@@
Rewards. The single-period reward for company $\mathcal{C}_{i}$ is solely based on its profit, reflecting the assumption that companies are profit-driven. The reward for investor $\mathcal{I}_{j}$ is the summation of two components: (1) the portfolio profit margin, and (2) the weighted average ESG score of the investor’s portfolio multiplied by the investor’s ESG preference, $\alpha^{\mathcal{I}_{j}}$.
@@
Social Outcome Metrics. We evaluate agent performance based on two key social outcome metrics: the final climate risk level, $P_{100}$, and the total market wealth at the end of the period, $W_{100}$. Additionally, we keep track of other social outcome metrics such as total number of severe climate events and number of companies that go bankrupt.
@@ 4.1 The trade-off for the agents frames climate change as a social dilemma
Companies’ trade-off between short-term private costs and long-term collective benefits create a social dilemma within the environment. To illustrate this, we use empirical Schelling diagrams (Hughes et al., 2018), which plot the focal-company payoffs from following either a cooperative or defecting policy, depending on the number of other cooperating company agents in a 5-company, 3-investor environment.
@@
Figure 3: Schelling diagrams demonstrating that the environment constitutes a social dilemma. The graphs compare payoffs between cooperation (mitigation, blue lines) and defection (no mitigation, red lines) for a focal company, given varying number of other cooperating companies. Yellow lines represent the average payoff across all companies when the focal company defects. Subfigure (a) illustrates the selfish scenario, where all three investors consistently prioritize financial returns ($\alpha^{\mathcal{I}_{j}}=0,\text{ for }j=1,2,3$). Here, defection always yields higher payoffs for the focal company than cooperation, leading all companies to defect. However, widespread defection results in lower overall profits, as the average payoff (yellow) increases with greater cooperation. Subfigure (b) and (c) correspond to two and three infinitely ESG-conscious investors ($\alpha^{\mathcal{I}_{j}}\approx\infty$), respectively. In (b), cooperation yields higher payoffs than defection for the focal company when few others cooperate. In (c), cooperation outperforms defection in all cases. Therefore, (b-c) demonstrate how investor behavior can transform the environment, eliminating the social dilemma by aligning corporate incentives with mitigation. Subfigures (d) and (e) build on (c) with three ESG-conscious investors. Subfigure (d) introduces resilience spending, while (e) adds greenwashing. The latter reintroduces a social dilemma, where corporations again avoid mitigation.
@@
In our case, cooperation represents altruistic investment in mitigating emissions. We simulate a cooperative policy as one which consistently invests 0.5% [^3] Since company actions are continuous, we selected 0.5% as a benchmark for illustrating the social dilemma structure in the Schelling diagram. This aligns with Patagonia’s cooperative approach, where $\sim$0.3of its valuation is allocated annually to climate initiatives. The diagram’s shape remains consistent across other tested values. of capital in mitigation and 0 in greenwashing or resilience building. In Figure 3(a) to 3(c), a defector policy takes zero action in all of mitigation, greenwashing, and resilience. As shown in Figure 3(a), when investors are solely profit-driven, the payoff for following a defector policy is always higher than the cooperator payoff. However, if all agents fail to cooperate, overall payoffs will be extremely low, showing that the environment represents a social dilemma. Figures 3(b) and 3(c) illustrate how the payoffs for cooperation and defection evolve as the proportion of ESG-conscious investors with positive ESG preferences increases. Notably, Figure 3(b) suggests the possibility of a bifurcated equilibrium when investor preferences are mixed. Specifically, climate-friendly companies may attract ESG-conscious investors, while other companies free ride and attract profit-motivated investors. Figure 3(c) shows that when all investors prioritize ESG-friendly firms, cooperation dominates defection in all cases. Therefore, with enough ESG-conscious investors the environment is no longer a social dilemma, and companies are actually greedily motivated to invest in mitigation to attract investors.
@@ 4.2 Multi-agent Reinforcement Learning Baselines
We are motivated to simulate companies and investors as independent, selfishly motivated agents that specialize in maximizing their own expected reward. We employ a state-of-the-art MARL algorithm based on Independent PPO (IPPO) (De Witt et al., 2020; Yu et al., 2022). so that each agent has its own policy parameters, and agents do not share parameters among themselves. At the back-end, InvestESG is implemented in both PyTorch (Paszke et al., 2017) and JAX (Bradbury et al., 2018), capable of simulating the investment behaviors in a massively parallel sense by vectorization on both CPUs and GPUs.
@@
By default, 5 companies are modeled with an initial capital of \$10 trillion each, alongside 3 investors, each starting with \$16 trillion. This creates a total beginning market wealth of \$98 trillion, which is comparable to the global stock market cap (WFE Statistics, 2023). We conduct the experiments in such a 5-company, 3-investor environment by default unless specified otherwise. All the following results are averaged result based on 3 runs of different random seeds, with the standard error plotted in the shaded region [^5] In simulating climate events as described in Equation equation 2, we use a fixed random seed across training episodes. This approach ensures that, while event occurrence varies with episode-specific climate risks, the underlying randomness remains deterministic. This setup is based on two considerations. First, unlike MARL agents learning from scratch, real-world companies possess a baseline understanding of trade-offs and can reasonably anticipate future risks; a deterministic training environment mirrors this insight. Second, our study aims not to develop agents that solve climate change under all stochastic scenarios, but rather to train agents that grasp the strategic trade-offs faced by companies and investors. A partially deterministic environment makes learning and convergence easier without compromising the study’s structural conclusions. For robustness, the environment setup allows agents to be trained in a fully stochastic environment to ensure generalizability. . Following economic literature that argues more than 10 agents are often needed to observe meaningful group size effects (Arifovic et al., 2023), we provide additional results in Appendix 11 from settings with 10 companies and 10 investors, as well as 25 companies and 25 investors, which are consistent with our findings with the 5-company, 3-investor setting. [^6] The 25-by-25 agent setting aligns with Zhang et al. (2022), which employs 27 agents in a MARL framework to address climate change issues.
5 Main Results
@@
Figure 4: Climate risk, total market wealth, total mitigation amount and climate event occurrence over the course of training for the 5-company-3-investor case. We compare the status quo scenario with solely profit-driven investors (investors with ESG consciousness level of 0), both with and without the ESG disclosure mandate, to scenarios involving three ESG-conscious investors with ESG consciousness level of $\alpha=0.5$, $\alpha=1$, and $\alpha=10$. These results indicate that merely disclosing ESG scores is insufficient to resolve the social dilemma if investors are not interested in investing in climate-friendly companies. Only when investors have a high level of ESG consciousness—$\alpha=10$, in this case—does the ESG mandate make a difference in increasing the level of mitigation.
Highly ESG-conscious investors are required to effectively incentivize mitigation efforts. To evaluate the effectiveness of ESG disclosure mandates in incentivizing emissions mitigation by companies, we conduct three experiments: (1) Status Quo: All agents are profit-motivated, and no ESG scores are released; (2) Status Quo with Mandate: ESG scores are disclosed, but investors remain profit-driven ($\alpha=0$); (3) Mandate with ESG-Conscious Investors: ESG scores are disclosed, and investors are ESG-conscious ($\alpha>0$). In all scenarios, companies only choose mitigation effort levels, with greenwashing and resilience spending disabled. Figure 4 tracks evaluation metrics over training. Initially, company agents act randomly, leading to frequent mitigation and optimal collective outcomes in the first 200 episodes. As training progresses, agents adopt profit-maximizing strategies, reducing mitigation efforts and achieving sub-optimal collective outcomes characteristic of social dilemmas. Over time, agents gradually realize that, in high-risk scenarios, investing in mitigation becomes marginally beneficial, leading to a slow increase in total market wealth.
The final system climate risks for the Status Quo with Mandate and low-ESG-conscious scenarios ($\alpha=0.5$ and $1$) are similar to those in the Status Quo, indicating that mandatory ESG disclosure alone does not significantly incentivize mitigation when investors prioritize profits. However, when investors have a high level of ESG-consciousness ($\alpha=10$), companies invest significantly in mitigation, reducing climate risk and increasing market wealth. In this setting, a few leading companies attract the majority of ESG-conscious investments, forming a positive feedback loop where better market performance draws further investments (see Appendix Figure 11(c) for a demonstration for such pattern in the 10 companies, 10 investors case).
These findings align with research showing that ESG disclosure mandates encourage climate-friendly efforts driven by societal and stakeholder pressures (Cormier and Magnan, 1999; Gamerschlag et al., 2011; Fama and French, 2007; Friedman and Heinle, 2016; Chen et al., 2018; Ioannou and Serafeim, 2019), and that fund managers are willing to sacrifice financial returns for ESG benefits (Krueger et al., 2021).
@@
Heterogeneous investor preferences lead to bifurcation in agent strategies. Research shows that investors have varying preferences for ESG efforts and respond in different ways (Amel-Zadeh and Serafeim, 2018). In this scenario, we initialized three investors with different levels of ESG-consciousness, with Investor 0 representing the solely profit-seeking investor with ESG-preference set to $\alpha^{\mathcal{I}_{0}}=0$, and Investor 1 and 2 representing highly ESG-conscious investors with $\alpha^{\mathcal{I}_{1}}=\alpha^{\mathcal{I}_{2}}=10$, which matches the scenario presented in the Schelling diagram in Figure 3(b). Figure 5(a) and 5(b) highlight a divergence in both investor and company behavior. Profit-driven investor 0 distributes investments evenly across companies, while ESG-focused investors (1 and 2) favor climate-conscious firms, such as Company 0, which prioritizes mitigation. Such divergence extends to company strategies: Company 0 learns to attract more ESG investment by focusing on mitigation, while others prioritize financial returns at the expense of some investor interest.
@@
The possibility of greenwashing does not significantly undermine mitigation efforts. Building on research that examines the existence and extent of greenwashing (Wu et al., 2020; Marquis et al., 2016; Siano et al., 2017), and the concern that the ESG disclosure policy can backfire with greenwashing (El-Hage, 2021), we explore whether companies adjust their strategies when greenwashing is permitted. Based on the Schelling diagram in Figure 3(d) and Figure 5(b), we predict that investors would be misled by greenwashing, prompting companies to prioritize greenwashing over mitigation due to its lower cost, leading to waste of resources. However, this prediction is not observed when simulating agents with IPPO. Figure 6(b) plots results under varying greenwashing costs, represented by coefficient $\beta$ in Equation 3, where a larger $\beta$ indicates cheaper greenwashing. All investors have an ESG-consciousness level of $\alpha=1$. The results show that regardless of greenwashing cost, companies initially explore greenwashing and mitigation equally but quickly drop greenwashing as the main strategy.
This likely occurs because, in the early episodes, investors have not yet linked ESG scores to their investment strategies, so do not respond to greenwashing efforts with increased investment. Instead, greenwashing presents an immediate cost for companies, which does not pay off with reduced climate risk, so they quickly learn to avoid it. In contrast, even without immediate investor rewards, companies may recognize the long-term benefits of mitigation (which leads to higher collective returns), and be incentivized to continue those efforts at first, even if they later learn to defect. This mirrors reality, where if investors are slow to adjust their strategies, companies may abandon greenwashing early but persist in low levels of mitigation for either its long-term climate benefits or the foresight of regulations and societal pressure that will eventually nudge them towards sustainable operation anyway. This is consistent with some empirical literature stating that despite the emergence of some greenwashing, ESG disclosure mandates overall encourage mitigation (Fiechter et al., 2022).
@@
Providing additional information about climate risk increases mitigation efforts. In this scenario, we want to observe the effect of having additional climate-related information available on mitigation behavior. Therefore, we provide the climate event probability and climate event occurrences as additional information in the observation space for both companies and investors. In the default 5-company-3-investor scenario, having additional climate-related information reduces the ending system climate risks, as shown in Figures 7(a). Figure 7(b) show similar effects of having additional climate-related information in the scenario where investors are not present. This result suggests that climate-related information helps both companies and investors make better environmental decisions. Even in the absence of ESG-focused investors, educating companies about the overall system improves outcomes. This aligns with literature showing that raising awareness encourages positive corporate responses (Delmas and Toffel, 2008; Bowen, 2000).